OpenStudy (anonymous):

How would I graph the Hyperbolas on a parent graph? y=1/x+5

OpenStudy (tkhunny):

Do you mean \(y = \dfrac{1}{x}+ 5\) as you have written or did you mean \(y = \dfrac{1}{x+5}\)?

OpenStudy (anonymous):

As I have written is what I meant! :)

OpenStudy (tkhunny):

Fine. A "Parent Graph" might be \(y = \dfrac{1}{x}\). Agreed?

OpenStudy (anonymous):

Yes and now I'm suppose to move it 5 units up. Which is +5 But how would I go about graphing it since it is a fraction

OpenStudy (tkhunny):

Why does the fraction make graphing more difficult? Just make a table or get a calculator to do it. Just don't pick x = 0.

OpenStudy (anonymous):

I don't understand how to graph this so yeah. Okay so if I don't pick x=0 I picked x=2 Then it'd be Y=1/2+5 which I would then proceed to graph up 1/2 and over 5?

OpenStudy (tkhunny):

For y = 1/x x y 1 1 2 1/2 3 1/3 4 1/4 -1 -1 -2 -1/2 -3 -1/3 -4 -1/4 For y = 1/x + 5 x y 1 1+5 = 6 2 1/2+5 = 11/2 3 1/3+5 = 16/3 4 1/4+5 = 21/4 -1 -1+5 = 4 -2 -1/2+5 = 9/2 -3 -1/3+5 = 14/3 -4 -1/4+5 = 19/4 Make sense?